OpenStudy (anonymous):

Solve the following system of equations. x - 2y = 6 2x - 4y = 10

OpenStudy (anonymous):

Wich one is the correct answer? a.Infinitely Many Solutions b. No Solutions c.(0, 0) d. (6, 10)

OpenStudy (anonymous):

I thik no solution because 2*(x-2y)=2*6 2x-4y=12 now 2x-4y=10 thus no solution

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

I'm sure